# Understanding Epistemic Language with a Bayesian Theory of Mind

### v1

@@ A.1 Experimental Procedure
The interface useds for collecting the statements and evaluating statements are shown in Figure A1. Participants first completed a tutorial that explained the task and experimental interface, then answered 5 comprehension questions before proceeding to the main experiment. In the main experiment, they were shown 10 out of the 20 stimuli in a randomized order.
@@ B.1 Belief-Space Sequential Inverse Plan Search (BSIPS)
Our BToM inference algorithm is a belief-space policy variant of Sequential Inverse Plan Search (SIPS, Zhi-Xuan et al. (2020)), which uses policies instead partial plans to evaluate action likelihoods as in more recent extensions of SIPS (Zhi-Xuan et al., 2024b, a). Perhaps surprisingly, Belief-Space SIPS (BSIPS) is able to exactly compute the posterior over beliefs, goals, and states (Equation 10) without any Monte Carlo approximation: For the scenarios we considered, there were between 120 and 5940 possible combinations of goals, initial states, and belief distributions, and enumerative inference over these hypothesis spaces could run as fast as 0.1s per action (120 hypotheses), going up to 20s per action (for 5490 hypotheses). Experiments were conducted with a i7-1370P 1.90 GHz CPU and 64 GB RAM.
Algorithm 1 provides the pseudocode for BSIPS. At each step $t$, we simulate how the environment changes, and how the agent updates their beliefs based on what they see (L6-7). Next, we efficiently compute belief-space $Q$-values by leveraging the $Q_{\text{MDP}}$ approximation described in Section 3.2.1. This involves averaging over $Q$-values for each state hypothesis $\tilde{s}$ in the agent’s belief $b_{t}^{i}$, which can be done cheaply by memoizing and reusing shortest-path computations across all belief hypotheses (L9-13). The $Q$-values allow us to compute the likelihood of the observed action $a_{t}$, allowing us to reweight each hypothesis by how well it explains the observations (L14-16).
For epistemic language evaluation at this scale, the technical challenges are mostly representational, not algorithmic. However, scaling to larger spaces of goals (Zhi-Xuan et al., 2024a) and (belief) states will require additional layers of Monte Carlo approximation. Our implementation in Gen (Cusumano-Towner et al., 2019) can naturally be extended to these cases, e.g. by leveraging Sequential Monte Carlo for approximate inference of initial states (Del Moral et al., 2006; Lew et al., 2023).
@@
B.2 Parameter Fitting
We performed a grid search over parameters for the LaBToM model. The range of inverse temperatures $\beta$ for the Boltzmann policy went from $0.5$ to $4$ in multiplicative increments of $\sqrt{2}$. This produced human correlations between $r=0.75$ (at $\beta=0.5$) and $r=0.81$ (at $\beta=4$) for current beliefs, and between $r=0.64$ (at $\beta=0.5$) and $r=0.80$ (at $\beta=2^{3/2}$), with $\beta=2^{3/2}$ producing the best fit overall.
We also fitted the threshold parameters $\Theta$ used in our ELoT representation. We performed grid-based coordinate ascent with a step size of 0.05, starting from values derived from the literature mapping modal words to probabilities Wesson and Pulford (2009); Hahn and Engelmann (2014); Meder et al. (2022), and limiting the search to a range of 0.2 above and below this starting point. Our initial and final threshold parameters are shown in Table B1. To evaluate threshold sensitivity, we also ran the same procedure to minimize correlation with humans. This produced a value of $r=0.71$, which was still much higher than the next best model (GPT-4o I+Na+FS, $r=0.52$).
@@ B.3 ELoT Translation Prompt
We used Gemini 1.5 Flash (Gemini Team et al., 2024) prompted with 34 examples to translate natural language statements into ELoT formulae. ELoT formulae were represented in a Prolog-like syntax analogous to the mathematical syntax we show in Table 1. The prompt used is shown below. We only show the first 10 out of 34 examples due to space constraints, and provide the full set in our code release.
@@ B.4 LLM Baseline Prompts
Appendix C Additional Results
C.1 Direct Translation w/o ELoT
To investigate the importance of our ELoT formalism, we also experimented with an ablated translation model which skips the ELoT representation and directly translates from natural language to the lower-level representation using Gemini 1.5 Flash with few-shot prompting. We find that this approach can handle most simple belief sentences, but makes errors when the statement becomes more complex, or involves distinguishing knowledge from belief. We highlight some failure cases in Table C2.
Example Translations and Errors
Statement: The player initially did not have confidence box 1 would hold the red key.
Trans.: prob_of(player, exists(and(key(K), iscolor(K, red)), inside(K, box1))) <= threshold(confidence)
Error: Non-existent threshold for confidence.
Statement: The player knows box 3 contains a blue key.
Trans.: prob_of(player, exists(and(key(K), iscolor(K, blue)), inside(K, box3))) >= threshold(knows)
Error: Non-existent threshold for knows, fails to assert that box 3 contains a blue key (knowledge should be factive).
Statement: The player is unsure which box has a key.
Trans.: forall(box(B), or(prob_of(player, exists(key(K), inside(K, B))) < threshold(uncertain), not(exists(key(K), inside(K, B)))))
Error: Incorrectly depends on whether box B actually has a key (uncertainty claims are not factive).
Table C2: Erroneous translations from English to probability comparisons, without using ELoT expressions.
C.2 Impact of a Normalized Statement Prior
@@
Table C3 shows the impact of making either assumption, in terms of the Pearson’s correlation coefficient (PCC) $r$ and mean absolute error (MAE) with human judgments. Consistent with Ying et al. (2024b), using a normalized statement prior largely improves the correlation while reducing the mean absolute error for the full LaBToM model. In other words, people appear more willing to say that a statement $\varphi$ is true only if they have evidence for $\varphi$, and otherwise default to a 50-50 rating. We see sharper differences for initial belief statements, which is likely because priors have a stronger effect on initial beliefs, whereas an agent’s current beliefs are more strongly determined by their percepts: If an agent sees that a box is empty, an observer’s judgment about whether the agent believes that the box is empty should not depend on the observer’s prior.
@@
C.3 Per-Factor Breakdown of Model Performance
@@
Table C4: Human vs. model correlations broken down by factors present in epistemic language.
Table C4 shows the correlation between human judgments and model outputs broken down by the annotated factors described in Table 2. LaBToM robustly outperforms the baselines on almost all of these data splits, achieving a correlation around $r=0.8$ in each case. The one exception is the set of statements about what the agent initially knows (Init. Know., $r=0.31$), such as “The player initially knew that the red key was in box 3”. As we discuss in the next section, this is likely because human participants assume that direct perception or justification is necessary for other agents to know some proposition $\phi$. In contrast, our model treats knowledge claims as equivalent to claims of true belief.
C.4 Differences in Human Ratings vs. Model Inferences
@@
People appear to adapt probability thresholds. Threshold uncertainty is closely related to another potential driver of difference: Unlike our model, participants appear to contextually adapt probability thresholds associated with modal words, in line with work on the pragmatics of epistemic modals (Schuster and Degen, 2020; Rudin, 2016; Lassiter, 2017). This is evinced by human responses for current belief statement S2 in Fig. 3b. People rate “The player is unsure which box has a key” highly despite the apparent confidence that the player exhibits in looking for a key in box 3 (at the expense of looking in box 1 or box 2). This is consistent with an upwards adjustment of $\theta_{\textsf{uncertain}}$ from 0.55, such that the player is judged as uncertain even when they seem to think it is quite likely for a key to be in box 3. In contrast, our model thinks it is unlikely that the player is uncertain. Similar effects can be seen for current statement S1 and S5 in the same scenario, except that people appear to adjust their threshold for may and might downwards from $0.3$ and $0.2$ to accommodate even highly unlikely possibilities.
@@
People reason boundedly about logical implications. A final difference between humans and our model is that people appear to exhibit bounded reasoning when evaluating statements with non-obvious implications. In Figure C2, for example, “The player initially believed a red key was most likely in box 2” is rated lowly by humans initially, then more highly, whereas our model assigns a zero rating to that statement by Judgment Point 2. This is because by then, it is clear that the player did not need a blue key, and was instead looking for a red key in box 1. The fact that they looked in box 1 first also implies that they initially thought box 1 was most likely to contain a red key, not box 2. Our model captures this reasoning, but people do not seem to independently grasp these multi-step implications, consistent with studies on the boundedness of human reasoning (Smets and Solaki, 2018; Mercier and Sperber, 2017).
@@
C.5 Misclassified In-Context vs. Out-of-Context Statements
As Table 4 in the main text shows, LaBToM accurately classifies most statements by assigning them a higher likelihood in their original context. However, about 30% of statements were assigned either equal or higher likelihoods out-of-context. By inspecting these cases more closely, we found that misclassification largely occurred due to the following reasons:
  • Dataset Noise. A few participants wrote statements that were implausible in their original context.
  • Similar Plausibility Across Contexts. Some statements described beliefs that were equally likely to be true out-of-context, or plausible in-context but even more plausible out-of-context.
  • Translation Errors. The ELoT translation step sometimes resulted in syntactically or semantically invalid translations. These statements were assigned a 0.5 likelihood across all contexts.
To illustrate the second point, we use an example statement from the scenario in Figure C2: “"The player believes that box 1 or 2 contains the red key.” By the end of the trajectory shown in Figure C2, this statement is assigned a normalized likelihood of 1.0 (since box 3 contains the blue key, the red key must be in box 1 or 2). In our two out-of-context trajectories, the player is shown finding a red key in box 1 and box 2 respectively. As such, it should be equally true that “"The player believes that box 1 or 2 contains the red key”, and our model assigns a likelihood of 1.0 in both cases. As a result, there is no difference in statement likelihoods in-context vs. out-of-context, even though our model is assigning statement likelihoods in an entirely reasonable manner.

### v2

@@ A.1 Experimental Procedure
The interface used for collecting the statements and evaluating statements are shown in Figure A1. Participants first completed a tutorial that explained the task and experimental interface, then answered 5 comprehension questions before proceeding to the main experiment. In the main experiment, they were shown 10 out of the 20 stimuli in a randomized order.
@@ B.1 Belief-Space Sequential Inverse Plan Search (BSIPS)
Our BToM inference algorithm is a belief-space policy variant of Sequential Inverse Plan Search (SIPS, Zhi-Xuan et al. (2020)), which uses policies to evaluate action likelihoods as in recent extensions of SIPS Zhi-Xuan et al. (2024b, a). Perhaps surprisingly, Belief-Space SIPS (BSIPS) is able to exactly compute the posterior over beliefs, goals, and states (Equation 11) without any Monte Carlo approximation: For the scenarios we considered, there were between 120 and 5940 possible combinations of goals, initial states, and beliefs. Enumerative inference over these hypotheses could run as fast as 0.1s/action (120 hypotheses), going up to 20s/action (5490 hypotheses). Experiments were conducted with a i7-1370P 1.90 GHz CPU and 64 GB RAM. Code can be found at https://osf.io/xq9c4/.
Algorithm 1 provides the pseudocode for BSIPS. At each step $t$, we simulate how the environment changes, and how the agent updates their beliefs based on what they see (L6-7). Next, we efficiently compute belief-space $Q$-values by leveraging the $Q_{\text{MDP}}$ approximation described in Section 3.2.1. This involves averaging over $Q$-values for each state hypothesis $\tilde{s}$ in the agent’s belief $b_{t}^{i}$, which can be done cheaply by memoizing and reusing shortest-path computations across all belief hypotheses (L9-13). The $Q$-values allow us to compute the likelihood of the observed action $a_{t}$, allowing us to reweight each hypothesis by how well it explains the observations (L14-16)
For epistemic language evaluation at this scale, the technical challenges are mostly representational, not algorithmic. However, scaling to larger spaces of goals Zhi-Xuan et al. (2024a) and (belief) states will require additional layers of Monte Carlo approximation. Our implementation in Gen Cusumano-Towner et al. (2019) can naturally be extended to these cases, e.g. by leveraging Sequential Monte Carlo for approximate inference of initial states Del Moral et al. (2006); Lew et al. (2023a).
@@
B.2 Parameter Fitting and Robustness Analyses
We performed a grid search over parameters for the LaBToM model. The range of inverse temperatures $\beta$ for the Boltzmann policy went from $0.5$ to $4$ in multiplicative increments of $\sqrt{2}$. This produced human correlations between $r=0.75$ (at $\beta=0.5$) and $r=0.81$ (at $\beta=4$) for current beliefs, and between $r=0.64$ (at $\beta=0.5$) and $r=0.80$ (at $\beta=2^{3/2}$), with $\beta=2^{3/2}$ producing the best fit overall. Across all values of $\beta$, we found that the full model outperformed the ablated baselines.
We also fitted the threshold parameters $\Theta$ used in our ELoT representation. We performed grid-based coordinate ascent with a step size of 0.05, starting from values derived from the literature mapping modal words to probabilities Wesson and Pulford (2009); Hahn and Engelmann (2014); Meder et al. (2022), and limiting the search to a range of 0.2 above and below this starting point. Our initial and final threshold parameters are shown in Table B1. To evaluate threshold sensitivity, we also ran the same procedure to minimize correlation with humans. This produced a value of $r=0.71$, which was still much higher than the next best model (GPT-4o I+Na+FS, $r=0.52$). Optimizing these thresholds for the True Belief and Non-Planning ablations led to maximal correlations of $r=0.12$ and $r=0.47$ respectively.
@@ B.3 ELoT Translation Prompt
For both grammar-constrained SMC and unconstrained sampling , we prompted the LLM with 36 examples to translate natural language statements into ELoT formulas. ELoT formulas were represented in a Prolog-like syntax analogous to the mathematical syntax we show in Table 1. The prompt used is shown below. We only show 10 out of 36 examples due to space constraints, and provide the full set in our code release. For translation of initial belief statements, we use a separate prompt to first translate the sentence from past tense to present tense, then translate the present tense sentence to ELoT.
@@ B.4 LLM Baseline Prompts
Appendix C Additional Results and Experiment Details
C.1 ELoT Parser Evaluation
[…]
LLaMa 3.1 8B + Grammar-Constrained SMC
Statement The player initially expected to find a key in box 3.
Gold Trans. believes(player, formula(exists(key(K), inside(K, box3))))
Translation
(67%) believes(player, likely(exists(key(K), inside(K, box3))))
(32%) believes(player, formula(exists(key(K), inside(K, box3))))
Error “Expected” is interpreted as “likely” instead of straightforward belief.
LLaMa 3.1 8B Unconstrained + Majority Vote
Statement The player initially thought that box 2 contained a red key.
Gold Trans. believes(player, formula(exists(and(key(K), iscolor(K, red)), inside(K, box2))))
Translation
(10%) believes(player, contains(red, box2))
(10%) type about(player, box2, contains)
Error Syntax errors, out of vocabulary term “contains”.
Gemini Flash 8B Unconstrained + Majority Vote
Statement The player initially did not know if box 2 had a key.
Gold Trans. not_knows_if(player, formula(exists(key(K), inside(K, box2))))
Translation (100%) not_knows_if(player, exists(key(K), inside(K, box2)))
Error Missing “formula” predicate to denote base formula.
Table C2: Erroneous ELoT translations for different parsing methods.
When evaluating translations, we performed both an exact equality check against our gold translations, and also a manual check for approximate semantic equivalence (i.e. cases where the parser gave a reasonable alternative translation of the original English statement). For replicability, we include these manual equivalence annotations in our code release.
Table C2 shows characteristic ELoT translation errors produced by each parsing method. Many of the the strict errors from grammar-constrained SMC pass the approximate equivalence check, and SMC often produces a distribution over reasonable alternative translations (Row 1). In contrast, unconstrained sampling from the same LLaMa 3.18B base model leads to many syntax errors that majority voting cannot correct for (Row 2). With Gemini Flash 8B, we observe the mode collapse issue that frequently afflicts finetuned LLMs O’Mahony et al. (2024), often leading all $n_{\sigma}$=$10$ samples to produce the same incorrect translation (Row 3).
C.2 Unimodal LLM Baseline with Symbolic PDDL Inputs
Prompt Format Current Beliefs Initial Beliefs
Image + Narrative + FS 0.52 (0.01) 0.41 (0.01)
PDDL + Narrative + FS 0.32 (0.01) 0.20 (0.01)
Table C3: GPT-4o correlation with human ratings across different input representations.
In our experiment, we used raw visual inputs, such as images or videos, for the multimodal LLM baselines. As the LaBToM model operates over symbolic representations of agent and environment states encoded in the Planning Domain Definition Language (PDDL), we also experimented with prompting LLMs with symbolic PDDL inputs. We report the results on the best performing LLM baseline (GPT4o I+Na+FS) in Table C3, which shows that replacing raw visual input with PDDL (as text) hinders performance.
C.3 Impact of a Normalized Statement Prior
@@
Table C4 shows the impact of making either assumption, in terms of the Pearson’s correlation coefficient (PCC) $r$ and mean absolute error (MAE) with human judgments, assuming the model has the gold translations. Consistent with Ying et al. (2024b), using a normalized statement prior largely improves the correlation while reducing the mean absolute error for the full LaBToM model. In other words, people appear more willing to say that a statement $\varphi$ is true only if they have evidence for $\varphi$, and otherwise default to a 50-50 rating. We see sharper differences for initial belief statements, which is likely because priors have a stronger effect on initial beliefs, whereas an agent’s current beliefs are more strongly determined by their percepts: If an agent sees that a box is empty, an observer’s judgment about whether the agent believes that the box is empty should not depend on the observer’s prior.
@@
C.4 Per-Factor Model Performance using Gold Translations
Table C5 shows the correlation between human judgments and model outputs when using the gold ELoT translations. Performance is broken down by the annotated factors described in Table 2. Using gold ELoT translations improves LaBToM’s correlation with human judgments from $r$=$0.76$ (Table 5) to $r$=$0.81$.
Additionally, LaBToM robustly outperforms the baselines on almost all of these data splits, achieving a correlation around $r$=$0.8$ in each case. The one exception is the set of statements about what the agent initially knows (Init. Know., $r=0.31$), such as “The player initially knew that the red key was in box 3”. As we discuss in the next section, this is likely because human participants assume that direct perception or justification is necessary for other agents to know some proposition $\phi$. In contrast, our model treats knowledge claims as equivalent to claims of true belief.
@@ C.4 Per-Factor Model Performance using Gold Translations
Table C5: Human vs. model correlations broken down by linguistic factors. LaBToM results use gold translations.
C.5 Differences in Human Ratings vs. Model Inferences
@@
People appear to adapt probability thresholds. Threshold uncertainty is closely related to another potential driver of difference: Unlike our model, participants appear to contextually adapt probability thresholds associated with modal words, in line with work on the pragmatics of epistemic modals Schuster and Degen (2020); Rudin (2016); Lassiter (2017). This is evinced by human responses for current belief statement S2 in Fig. 3b. People rate “The player is unsure which box has a key” highly despite the apparent confidence that the player exhibits in looking for a key in box 3 (at the expense of looking in box 1 or box 2). This is consistent with an upwards adjustment of $\theta_{\textsf{uncertain}}$ from 0.55, such that the player is judged as uncertain even when they seem to think it is quite likely for a key to be in box 3. In contrast, our model thinks it is unlikely that the player is uncertain. Similar effects can be seen for current statement S1 and S5 in the same scenario, except that people appear to adjust their threshold for may and might downwards from $0.3$ and $0.2$ to accommodate even highly unlikely possibilities.
@@
People reason boundedly about logical implications. A final difference between humans and our model is that people appear to exhibit bounded reasoning when evaluating statements with non-obvious implications. In Figure C2, for example, “The player initially believed a red key was most likely in box 2” is rated lowly by humans initially, then more highly, whereas our model assigns a zero rating to that statement by Judgment Point 2. This is because by then, it is clear that the player did not need a blue key, and was instead looking for a red key in box 1. The fact that they looked in box 1 first also implies that they initially thought box 1 was most likely to contain a red key, not box 2. Our model captures this reasoning, but people do not seem to independently grasp these multi-step implications, consistent with studies on the boundedness of human reasoning Smets and Solaki (2018); Mercier and Sperber (2017).